Find the solutions to 6x^2 + 5x - 4= 0. Is the sum of your solutions equal to
-b/a? Explain why the sum of the solutions to any quadratic equation is -b/a?.
(Hint: Use the quadratic formula.)

Answer by stanbon(75887)

Find the solutions to 6x^2 + 5x - 4= 0. Is the sum of your solutions equal to
-b/a?
Explain why the sum of the solutions to any quadratic equation is -b/a?.
(Hint: Use the quadratic formula.)
---
1st solution: x1 = (-b/2a)+(sqrt(b^2-4ac)/(2a)
---
2nd solution: x2 = (-b/2a)-(sqrt(b^2-4ac)/(2a)
---
Adding you get: x1+x2 = 2(-b/2a) = -b/a
============================================
Cheers,
Stan H.